OpenStudy (anonymous):

Find the area between the curves. y=sin2x, y=cos2x, x=0 Please help!!!

OpenStudy (dumbcow):

\[A = \int\limits_{0}^{\pi/8}\cos(2x) -\sin(2x) dx\]